# MCGrad: Multicalibration at Web Scale

### v2

@@ A.2. Convergence Analysis
The gradient of the loss function with respect to $w_{k}$ is equal to
@@
where $f_{t+1}(x)=\mu(F_{t}(x)+\sum_{k=1}^{K}w_{k}h_{k}(x,f_{t}(x)))$. A GBM can be shown to correspond to a coordinate gradient descent in the coefficient space (Lu and Mazumder, 2020), defined by
$w^{m+1}=w^{m}-\rho\nabla_{j_{m}}\mathcal{L}_{t}(w^{m})e_{j_{m}}$
where $e_{j}$ is the $j$-th basis vector in $\mathbb{R}^{K}$ and $\rho$ is a step size.
A.2.1. Decreasing Loss Property
Proposition A.0.
Assume that the loss function $\mathcal{L}$ is a convex, $L$-smooth loss function (for log loss, $L=1/4$) and the step size $\rho$ is set as $\rho=1/L$. Then, for every $t\geq 0$,
@@ A.2.1. Decreasing Loss Property
where the inequality is strict whenever $\nabla_{j_{m}}\mathcal{L}_{t}(w^{m})\neq 0$ in some iteration $m\in\{0,1,\ldots,M_{t}-1\}$.
@@
First, note the following relations, for any $\rho_{m}$:
$\mathcal{L}_{t}(w^{m+1})$ $\leq$ $\mathcal{L}_{t}(w^{m}+\rho_{m}e_{j_{m}})$
$\leq$ $\mathcal{L}_{t}(w^{m})+\rho_{m}\nabla\mathcal{L}_{t}(w^{m})^{\top}e_{j_{m}}+\frac{L}{2}\rho_{m}^{2}.$
@@
It follows that
$\mathcal{L}_{t}(w^{M_{t}})-\mathcal{L}_{t}(w^{0})\leq-\frac{1}{2L}\sum_{m=0}^{M_{t}-1}(\nabla_{j_{m}}\mathcal{L}_{t}(w^{m}))^{2}.$
Hence, it follows that $\mathcal{L}_{t}(w^{M_{t}})-\mathcal{L}_{t}(w^{0})\leq 0$ where the inequality is strict whenever $\nabla_{j_{m}}\mathcal{L}_{t}(w^{m})\neq 0$ for some iteration $m\in\{0,1,\ldots,M_{t}-1\}$. ∎
For a GBM that in each iteration greedily selects a weak learner with the largest absolute value of the loss-gradient, i.e.,
$j_{m}\in\arg\max_{j}|\nabla_{j}\mathcal{L}_{t}(w^{m})|,$
we can establish a rate of convergence to the loss of the optimum linear combination of weak learners, using Theorem 4.2 (Lu and Mazumder, 2020). Let $w^{*}(t)=\arg\min_{w}\mathcal{L}_{t}(w)$. Then,
@@
where $\mathrm{Dist}_{t}$ is a distance between the level set $\{w:\mathcal{L}_{t}(w)\leq\mathcal{L}_{t}(w^{0})\}$ and $w^{*}(t)$, given in (Lu and Mazumder, 2020). This provides a theoretical justification for assuming updating the predictor with an optimal ensemble of weak learners, minimizing the loss function, in each round of the algorithm.
A.2.2. A Bound on the Multicalibration Error
We next provide a bound on the multicalibration error of the predictor $f_{T}$.
Proposition A.0.
For any $T>0$, $f_{T}$ is $\alpha$-MC with
$\alpha\leq\frac{2}{\sqrt{3}}C_{T}\sqrt{2\ \mathbb{E}\left[|F_{T+1}(X)-F_{T}(X)|\right]+\frac{1}{8}\mathbb{E}\left[(F_{T+1}(X)-F_{T}(X))^{2}\right]+\epsilon_{T}}$
where
$C_{T}^{2}:=\max_{h\in\mathcal{H},g\in\mathcal{G}}\frac{\mathbb{E}\left[g(X)\mid h(X)=1\right]}{\mathbb{E}\left[f_{T}(X)(1-f_{T}(X))\mid h(X)=1\right]}$
and $\epsilon_{T}:=\mathbb{E}\left[\mathcal{L}(F_{T+1}(X),Y)\right]-\mathcal{L}_{t}(w^{*}(T))$.
Before presenting the proof of the proposition, we discuss the result of the proposition.
The proposition provides a bound on the miscalibration error in terms of the gap between successive predictor values $F_{T+1}$ and $F_{T}$, $\epsilon_{T}$ that captures the gap between the loss function values achieved by the ensemble of regression trees added by the GBM and the optimal ensemble in round $T$, and $C_{T}$ that depends on how much the predictor $f_{T}$ concentrates around the boundary values 0 and $1$. The dependence on the gap between $F_{T}$ and $F_{T+1}$ is such that the multicalibration error decreases as the gap between $F_{T}$ and $F_{T+1}$ decreases. The error term $\epsilon_{T}$ can be controlled by the number of iterations $M_{t}$ used by GBM in round $t$ of the algorithm, as demonstrated in Equation (5).
For the $C_{T}$ term, note:
This bound is finite if and only if $\mathbb{E}\left[f_{T}(X)(1-f_{T}(X))\mid h(X)=1\right]>0$ for every group $h\in\mathcal{H}$. In other words, for every group $h$, there exists a member $x$ such that $0<f_{T}(x)<1$. This condition clearly holds if $0<f_{T}(x)<1$ for every $x\in\mathcal{X}$ which holds whenever $F_{T}(x)\in[-R,R]$ for every $x\in\mathcal{X}$, for some constant $R>0$. If the predictor $f_{T}$ takes values in $\epsilon,1-\epsilon]$, for some $0<\epsilon\leq 1/2$, then clearly $C_{T}^{2}\leq 2/\epsilon$.
@@ A.2.2. A Bound on the Multicalibration Error
We first show the following lemma:
Lemma A.4.
If $f_{T}$ is such that $|\mathbb{E}\left[h(X)g(f_{T}(X))(y-f_{T}(X))\right]|\geq\alpha\sqrt{\mathbb{E}\left[h(X)f_{T}(X)(1-f_{T}(X))\right]}$, for some $h\in\mathcal{H}$ and $g\in\mathcal{G}$, then there exists $\tilde{w}\in\mathbb{R}^{K}$ such that for $\tilde{h}(x)=\sum_{k=1}^{K}\tilde{w}_{k}h_{k}(x,f_{t}(x))$, we have
where
$C_{T}^{2}:=\max_{h\in\mathcal{H},g\in\mathcal{G}}\frac{\mathbb{E}\left[g(X)\mid h(X)=1\right]}{\mathbb{E}\left[f_{T}(X)(1-f_{T}(X))\mid h(X)=1\right]}.$
Proof.
For any convex, $L$-smooth loss function $\mathcal{L}$, and any $F$ and $h$, we have
$\mathcal{L}(F(x),y)-\mathcal{L}(F(x)+h(x),y)$
$\geq$ $\frac{d}{dF}\mathcal{L}(F(x)+h(x),y)(-h(x))$
@@
For the log loss $(d/dF)\mathcal{L}(F,y)=-(y-\sigma(F))$, hence, we have
$\mathbb{E}\left[\mathcal{L}(F(X),Y)\right]-\mathbb{E}\left[\mathcal{L}(F(X)+h(X),Y)\right]$
$\geq$ $\mathbb{E}\left[(Y-f(X))h(X)\right]-L\mathbb{E}\left[h(X)^{2}\right].$
Let $\tilde{w}$ be defined as:
@@
if $(h,g)=(h^{*},g^{*})$ and $\tilde{\theta}_{h,g}=0$, otherwise, where
$(h^{*},g^{*})\in\arg\max_{h\in\mathcal{G},g\in\mathcal{G}}\frac{|\mathbb{E}\left[h(X)g(f_{t}(X))(Y-f_{t}(X))\right]|}{\sqrt{\mathbb{E}\left[h(X)f_{t}(X)(1-f_{t}(X))\right]}}.$
For $w=\tilde{w}$, we have
@@
$\mathbb{E}\left[\mathcal{L}(F_{t}(X),Y)\right]-\mathbb{E}\left[\mathcal{L}\left(F_{t}(X)+\sum_{k=1}^{K}\tilde{w}_{k}h_{k}(X,f_{t}(X)),Y\right)\right]$
@@
$\geq$ $\left(1-L\right)\frac{\mathbb{E}\left[h^{*}(X)g^{*}(f_{t}(X))(Y-f_{t}(X))\right]^{2}}{\mathbb{E}\left[h^{*}(X)^{2}\right]}$
$=$ $\left(1-L\right)\frac{\mathbb{E}\left[h^{*}(X)f_{t}(X)(1-f_{t}(X))\right]}{\mathbb{E}\left[h^{*}(X)g^{*}(f_{t}(X))\right]}\cdot$
$\cdot\left(\frac{\mathbb{E}\left[h^{*}(X)g^{*}(f_{t}(X))(Y-f_{t}(X))\right]}{\sqrt{\mathbb{E}\left[h^{*}(X)f_{t}(X)(1-f_{t}(X)\right]}}\right)^{2}$
@@
Using the fact that for the log loss function $L=1/4$ and the definition of $C_{T}$, we have
@@
which completes the proof of the lemma. ∎
@@
$\mathbb{E}\left[\mathcal{L}(F_{t}(X),Y)\right]-\mathbb{E}\left[\mathcal{L}\left(F_{t}(X)+\sum_{k=1}^{K}w_{k}h_{k}(X,f_{t}(X)),Y\right)\right]$
$\leq$ $\mathbb{E}\left[\mathcal{L}(F_{t}(X),Y)\right]-\mathcal{L}_{t}(w^{*}(t))$
$=$ $\mathbb{E}\left[\mathcal{L}(F_{t}(X),Y)\right]-\mathbb{E}\left[\mathcal{L}(F_{t+1}(X),Y)\right]$
$+\mathbb{E}\left[\mathcal{L}(F_{t+1}(X),Y)\right]-\mathcal{L}_{t}(w^{*}(t)).$
We next consider the two terms, $\mathbb{E}\left[\mathcal{L}(F_{t}(X),Y)\right]-\mathbb{E}\left[\mathcal{L}(F_{t+1}(X),Y)\right]$ and $\mathbb{E}\left[\mathcal{L}(F_{t+1}(X),Y)\right]-\mathcal{L}_{t}(w^{*}(t))$. The first term captures the gap of the loss function values at two successive rounds of the algorithm. This term can be bounded as follows:
$\mathbb{E}\left[\mathcal{L}(F_{t}(X),Y)\right]-\mathbb{E}\left[\mathcal{L}(F_{t+1}(X),Y)\right]$
$\leq$ $\mathbb{E}\left[\frac{d}{dF}\mathcal{L}(F_{t+1}(X),Y)\cdot(F_{t}(X)-F_{t+1}(X))+\frac{L}{2}(F_{t}(X)-F_{t+1}(X))^{2}\right]$
$=$ $\mathbb{E}\left[(y-f_{t+1})(F_{t}(X)-F_{t+1}(X))+\frac{L}{2}(F_{t}(X)-F_{t+1}(X))^{2}\right]$
$\leq$ $\mathbb{E}\left[|Y-f_{t+1}(X)||F_{t}(X)-F_{t+1}(X)|+\frac{L}{2}(F_{t}(X)+F_{t+1}(X))^{2}\right]$
$\leq$ $2\ \mathbb{E}\left[|F_{t+1}(X)-F_{t}(X|\right]+\frac{L}{2}\mathbb{E}\left[(F_{t+1}(X)-F_{t}(X))^{2}\right].$
@@
The second term, $\mathbb{E}\left[\mathcal{L}(F_{t+1}(X),Y)\right]-\mathcal{L}_{t}(w^{*}(t))$ is the gap between the loss function value at the end of the $t$-th round of the algorithm and the loss function value obtained by adding an optimal linear combination of weak classifiers. This term can be controlled by the number of iterations used by GBM in round $t$. Specifically, this term can be bounded as given in Equation (5).
The assertion of the proposition follows by combining the above relations with Lemma A.4. ∎
Appendix B Benchmark Datasets
We include in our experimental analysis a total of $11$ datasets. Specifically:
  • American Community Survey (Ding et al., 2021): We use the folktable python package (Flood et al., 2021) to access the data and define six binarized prediction tasks (i.e., create six datasets): income (ACSIncome), mobility (ACSMobility), employment (ACSEmploy), travel time (ACSTravel), public health insurance (ACSPublic), health insurance (ACSHealth) . Also, ACSIncomeCA considers only the rows with the country set to California. For each target we use the features specified in the folktable package and add all features used in the specifications protected groups in Hansen et al. (2024);
  • UCI Bank Marketing: The dataset Marketing has the task of predicting if a client will subscribe to a term deposit, using features like housing, occupation, education, and age (Moro et al., 2014);
  • UCI Default of Credit Card Clients: The dataset Credit has the task of predicting credit card debt default using payment history and identity attributes (Yeh, 2009);
  • Home Mortgage Disclosure Act: The dataset HMDA has the task of predicting the acceptance of mortgage applications in 2017, using identity attributes of associated applicants (Cooper et al., 2023);
  • Medical Expenditure Panel Survey: The dataset MEPS has the task of predicting whether a household makes at least $10$ medical visits, based on socioeconomic and geographic information (Sharma et al., 2021).
Table 4 shows the number of samples, features and the class prior of each dataset.
Dataset # Samples $n$ # Features $d$ Class Distr.
MEPS 11079 139 0.169
Credit 30000 118 0.221
Marketing 45211 41 0.116
HMDA 114185 89 0.752
ACSIncomeCA 195665 10 0.410
ACSMobility 616207 23 0.735
ACSPublic 1123374 24 0.293
ACSTravel 1458542 19 0.437
ACSIncome 1655429 10 0.370
ACSEmploy 3207990 20 0.456
ACSHealth 3207990 31 0.150
Table 4. Number of samples, features and class distribution ($p(Y=1)$) for each dataset used in the benchmark evaluation.
Appendix C Setting Default Hyperparameters for MCGrad
MCGrad uses LightGBM as GBDT implementation. We found that order to get ML engineers to adopt MCGrad into their ML production system, is important that it works out-of-the-box. That is, it should work for their ML model, without application-specific tuning of the LightGBM hyperparameters. Making it possible for ML engineers to use MCGrad without tuning LightGBM hyperparameters substantially reduces the barrier to use MCGrad, and increases trust that it can be safely incorporated into automated ML model training pipelines.
In order to set default hyperparameters for MCGrad, we performed a grid search of LightGBM hyperparameters where we evaluated each hyperparameter configuration on a collection of 35 company-internal datasets. The purpose of this analysis was to select a set of LightGBM parameters that are suitable as default hyperparameters in MCGrad. We selected the hyperparameter configuration that across those 35 datasets (1) never degraded log loss and Area Under the Precision-Recall Curve (AUPRC) relative to the model scores before calibration adjustment, and (2) minimised the average MultiCalibration Error (MCE) across the datasets.
This resulted in the following set of hyperparameters:
learning_rate:
$0.02873$
max_depth:
55
min_child_samples:
$160$
n_estimators:
$94$
num_leaves:
55
lambda_l2:
$0.00913$
min_gain_to_split:
$0.15$
All other hyperparameters were left to LightGBM default values.

### v3

@@ Appendix B Convergence Analysis
The gradient of the loss function with respect to $w_{k}$ is given by
@@
where $f_{t+1}(x)=\mu(F_{t}(x)+\sum_{k=1}^{K}w_{k}h_{k}(x,f_{t}(x)))$.
A GBM can be shown to correspond to a coordinate gradient descent in the coefficient space (Lu and Mazumder, 2020), with updates of the form $w^{m+1}=w^{m}-\rho\nabla_{j_{m}}\mathcal{L}_{t}(w^{m})e_{j_{m}}$, where $e_{j}$ is the $j$-th standard basis vector in $\mathbb{R}^{K}$ and $\rho$ is the step size.
Decreasing Loss Property
Proposition B.0.
Assume that the loss function $\mathcal{L}$ is a convex and $L$-smooth (for the log loss, $L=1/4$) and that the step size $\rho$ is set as $\rho=1/L$. Then, for every $t\geq 0$, the following inequality holds
@@
[…]
@@
Note the following relations, for any $\rho_{m}$:
$\mathcal{L}_{t}(w^{m+1})\!\leq\!\mathcal{L}_{t}(w^{m}+\rho_{m}e_{j_{m}})\!\leq\!\mathcal{L}_{t}(w^{m})+\rho_{m}\nabla\mathcal{L}_{t}(w^{m})^{\top}e_{j_{m}}\!+\frac{L}{2}\rho_{m}^{2}.$
@@
It follows that $\mathcal{L}_{t}(w^{M_{t}})-\mathcal{L}_{t}(w^{0})\leq-\frac{1}{2L}\sum_{m=0}^{M_{t}-1}(\nabla_{j_{m}}\mathcal{L}_{t}(w^{m}))^{2}$, which allows us to conclude that $\mathcal{L}_{t}(w^{M_{t}})-\mathcal{L}_{t}(w^{0})\leq 0$ where the inequality is strict if $\nabla_{j_{m}}\mathcal{L}_{t}(w^{m})\neq 0$ for some $0\leq m<M_{t}$. ∎
For a GBM that, in each iteration, greedily selects the weak learner with the largest absolute value of the loss gradient, i.e., $j_{m}\in\arg\max_{j}|\nabla_{j}\mathcal{L}_{t}(w^{m})|,$ we can establish a convergence rate to the loss of the optimum linear combination of weak learners using Theorem 4.2 (Lu and Mazumder, 2020). Let $w^{*}(t)=\arg\min_{w}\mathcal{L}_{t}(w)$. Then,
@@
where $\mathrm{Dist}_{t}$ denotes the distance between the level set $\{w:\mathcal{L}_{t}(w)\leq\mathcal{L}_{t}(w^{0})\}$ and $w^{*}(t)$, as defined in (Lu and Mazumder, 2020).
This result justifies modeling each round as updating the predictor with an optimal ensemble of weak learners that minimizes loss. Proposition B.1 shows the training loss never increases with more rounds and converges to a limit as rounds go to infinity.
A Bound on the Multicalibration Error.
The bound in Proposition B.3 provides a theoretical upper-bound guarantee, showing that the multicalibration error converges to zero when $C_{T}$ is bounded by a constant, the gap between prediction values $F_{T+1}$ and $F_{T}$ vanishes as the number of rounds $T$ increases, and the GBM achieves optimal loss in round $T$. Before proving this, we need a preliminary result:
Lemma B.2.
If $f_{T}$ is such that $|\mathbb{E}\left[h(X)g(f_{T}(X))(y-f_{T}(X))\right]|\geq\alpha\sqrt{\mathbb{E}\left[h(X)f_{T}(X)(1-f_{T}(X))\right]}$, for some $h\in\mathcal{H}$ and $g\in\mathcal{G}$, then there exists $\tilde{w}\in\mathbb{R}^{K}$ such that for $\tilde{h}(x)=\sum_{k=1}^{K}\tilde{w}_{k}h_{k}(x,f_{T}(x))$:
$\text{ where }\ C_{T}^{2}\coloneqq\max_{h\in\mathcal{H},g\in\mathcal{G}}\frac{\mathbb{E}\left[g(f_{T}(X))\mid h(X)=1\right]}{\mathbb{E}\left[f_{T}(X)(1-f_{T}(X))\mid h(X)=1\right]}.$
@@
For any $F$, $h$ and any convex, $L$-smooth loss function $\mathcal{L}$,
$\mathcal{L}(F(x),y)-\mathcal{L}(F(x)+h(x),y)\geq\frac{d}{dF}\mathcal{L}(F(x)+h(x),y)(-h(x))$
@@
For the log loss $(d/dF)\mathcal{L}(F,y)=-(y-\mu(F))$, hence, we have
$\mathbb{E}\left[\mathcal{L}(F(X),Y)\!-\!\mathcal{L}(F(X)\!+\!h(X),Y)\right]\!\geq\!\mathbb{E}\left[(Y\!-\!f(X))h(X)\right]\!-\!L\mathbb{E}\left[h(X)^{2}\right]$
Let
$(h^{*},g^{*})\in\arg\max_{h\in\mathcal{G},g\in\mathcal{G}}\frac{|\mathbb{E}\left[h(X)g(f_{T}(X))(Y-f_{T}(X))\right]|}{\sqrt{\mathbb{E}\left[h(X)f_{T}(X)(1-f_{T}(X))\right]}}$
and, let $\tilde{w}$ be defined as:
@@
if $(h,g)=(h^{*},g^{*})$, otherwise $\tilde{w}_{h,g}=0$. For $w=\tilde{w}$, we have both
@@
$\mathbb{E}\left[\mathcal{L}(F_{T}(X),Y)\!-\!\mathcal{L}\left(F_{T}(X)\!+\!\!\sum_{k=1}^{K}\tilde{w}_{k}h_{k}(X,f_{T}(X)),\!Y\right)\right]$
@@
Because $L=1/4$ for the log loss, we conclude the proof
@@
Now, we can derive the upper bound of the MC error:
Proposition B.0.
$f_{T}$ is $\alpha$-MC with $\alpha\leq\frac{2}{\sqrt{3}}C_{T}\sqrt{\phi_{T}+\epsilon_{T}}$ where
$\phi_{T}$ $\coloneqq$ $2\ \mathbb{E}\left[|F_{T+1}(X)-F_{T}(X)|\right]+\frac{1}{8}\mathbb{E}\left[(F_{T+1}(X)-F_{T}(X))^{2}\right]$
$\epsilon_{T}$ $\coloneqq$ $\mathbb{E}\left[\mathcal{L}(F_{T+1}(X),Y)\right]-\mathcal{L}_{T}(w^{*}(T)),$
and $C_{T}$ is defined as in Lemma B.2.
Proof.
@@
$\mathbb{E}\left[\mathcal{L}(F_{t}(X),Y)\right]-\mathbb{E}\left[\mathcal{L}\left(F_{t}(X)+\sum_{k=1}^{K}w_{k}h_{k}(X,f_{t}(X)),\!Y\right)\right]\qquad$
$\leq\underbrace{\!\mathbb{E}\left[\!\mathcal{L}(F_{t}(X),\!Y)\!-\!\mathcal{L}(F_{t+1}(X),\!Y)\right]}_{(*)}\!+\!\underbrace{\mathbb{E}\left[\!\mathcal{L}(F_{t+1}(X),\!Y)\right]\!-\!\mathcal{L}_{t}(w^{*}(t))}_{(**)}$
@@
The first term (*) captures the gap of the loss function values at two successive rounds of the algorithm and can be bounded by $2\mathbb{E}\left[|F_{t+1}(X)-F_{t}(X|\right]+\frac{L}{2}\mathbb{E}\left[(F_{t+1}(X)-F_{t}(X))^{2}\right]$. The second term (**) measures the gap between the loss at the end of round $t$ and the loss from an optimal linear combination of weak classifiers. This gap can be reduced by increasing the number of GBM iterations and is bounded as in Eq (4). The lemma follows from Lemma B.2. ∎
The bound on the multicalibration error depends on three terms: (a) $C_{T}$, which depends inversely on the average prediction variance within subgroups, (b) $\phi_{T}$, which depends on the gap between the prediction values $F_{T+1}$ and $F_{T}$, and (c) $\epsilon_{T}$, which measures the gap between the loss achieved by the predictor $F_{T+1}$ and the optimum loss in round $T$. If $C_{T}$ is bounded by a constant, the MC error can be made arbitrarily small by reducing $\phi_{T}$ and $\epsilon_{T}$. As $T$ increases, $\phi_{T}$ approaches zero, and $\epsilon_{T}$ can be minimized by increasing the number of GBM iterations $M_{T}$, as shown in Eq. (4). Also, note that
This bound is finite if and only if $\mathbb{E}\left[f_{T}(X)(1-f_{T}(X))\mid h(X)=1\right]>0$ for every group $h\in\mathcal{H}$. Equivalently, for every group $h$, there exists at least one member $x$ such that $0<f_{T}(x)<1$. This condition clearly holds if $0<f_{T}(x)<1$ for all $x\in\mathcal{X}$ which is satisfied whenever $F_{T}(x)\in[-R,R]$ for every $x\in\mathcal{X}$, for some constant $R>0$. Moreover, if the predictor $f_{T}$ takes values in $\epsilon,1-\epsilon]$, for some $0<\epsilon\leq 1/2$, then $C_{T}^{2}\leq\frac{2}{\epsilon}.$
Appendix C Extended Experimental Results